OpenStudy (anonymous):

Elimination method: 4x – 5y = -31 -12x + 15y = 93

OpenStudy (anonymous):

multiply first equation by 3 and add bouth equations, you will get 0=0 It means bouth equations are the same. Infinitly many solutions

OpenStudy (lgbasallote):

i have a feeling this is no solution

OpenStudy (anonymous):

Times the top equation by 3. 12x-15y=-93 +-12x+15y=93 -------------- 0x+0y=0

OpenStudy (anonymous):

(1) 4x – 5y = -31 (2) -12x + 15y = 93 (1) 3(4x-5y=31) 12x-15y=93 -12x+15y=-93 x=0 y=0 This what I got but was not sure of it. So the answer would be no answer or infinate

OpenStudy (anonymous):

infinite

OpenStudy (anonymous):

12x-15y=93 -12x+15y=-93 --------------- 0x+0y = 0 so for any x and any y this will be correct. It means infinitly many x and y are solutions

OpenStudy (anonymous):

Thank you so much.